OpenStudy (tinylittlehelper):

Describe a rectangle whose perimeter is 132 ft. and whose area is less than 1 ft ^ 2. Is it possible to find a circle whose circumference is 132 ft. and whose area is less than 1 ft ^ 2? If not, provide an example or write a sentence explaining why no such circle exists. Please list out steps; I suck at this ;p

jhonyy9 (jhonyy9):

for a circle : perimeter P = 2*pi*r using this above wrote P = 132 so 132= 2*pi*r -divide both sides by 2 66 = pi*r - so use for pi = 3,14 and divide both sides by 3,14 66/3,14 = r r = 21,019 let be 21 r=21 area of circle = pi*r^2 area the above circle = 3,14*21^2 = 3,14*441 - so this not will be never equal 1 ft^2 hope these will can clarify you in this case of this exercise why not can being possibly true this above wrote ...

jhonyy9 (jhonyy9):

than are you here - so do you understand it now clearly sure ?

OpenStudy (tinylittlehelper):

I understand the circle part clearly now. Thank you!

jhonyy9 (jhonyy9):

np was my pleasure - good luck bye